### v1

# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

Abstract
Music recordings often suffer from audio quality issues such as excessive reverberation, distortion, clipping, tonal imbalances, and a narrowed stereo image, especially when created in non-professional settings without specialized equipment or expertise. These problems are typically corrected using separate specialized tools and manual adjustments. In this paper, we introduce SonicMaster, the first unified generative model for music restoration and mastering that addresses a broad spectrum of audio artifacts with text-based control. SonicMaster is conditioned on natural language instructions to apply targeted enhancements, or can operate in an automatic mode for general restoration. To train this model, we construct the SonicMaster dataset, a large dataset of paired degraded and high-quality tracks by simulating common degradation types with nineteen degradation functions belonging to five enhancements groups: equalization, dynamics, reverb, amplitude, and stereo. Our approach leverages a flow-matching generative training paradigm to learn an audio transformation that maps degraded inputs to their cleaned, mastered versions guided by text prompts. Objective audio quality metrics demonstrate that SonicMaster significantly improves sound quality across all artifact categories. Furthermore, subjective listening tests confirm that listeners prefer SonicMaster’s enhanced outputs over the original degraded audio, highlighting the effectiveness of our unified approach. Samples, code, model, dataset, and demo are all available through https://amaai-lab.github.io/SonicMaster/
Introduction
Music tracks produced in amateur settings often suffer from a variety of audio quality issues that distinguish them from professionally mastered recordings (Wilson and Fazenda 2016; Mourgela et al. 2024; Deruty and Tardieu 2014). For instance, an enthusiast recording vocals in a garage may introduce excessive reverberation, making the voice sound distant and “echoey.” Similarly, using inexpensive microphones or misconfigured interfaces can lead to distortion and clipping when loud peaks exceed the recording range, resulting in harsh crackles or flattened dynamics (Zang et al. 2025). Tonal imbalances are also common: a home recording might sound overly “muddy” or “tinny” if certain frequency bands dominate or vanish due to poor room acoustics or improper microphone placement. Even the stereo image can be narrowed or skewed, reducing the sense of space in the mix. In practice, engineers address these problems with separate specialized tools: e.g., dereverberation plugins to remove room echo, declipping algorithms to reconstruct saturated peaks, equalizers to rebalance frequencies, and stereo enhancers to widen the image. Mastering a flawed track thus becomes a labor-intensive process requiring expert skill and multiple stages of manual adjustment.
The need for an automated all-in-one solution is evident. Creators with limited resources often lack the expertise to apply the right combination of restoration tools, and a piecemeal approach may fail to fully recover a track’s fidelity. This motivates SonicMaster, a unified approach to music restoration and mastering that can correct a broad spectrum of audio degradations within a single model.
Figure 1: SonicMaster pipeline: a flow-matching model trained on the SonicMaster dataset (19 simulated enhancements across EQ, dynamics, reverb, amplitude, stereo) takes an input music waveform and text prompts to produce a high-fidelity enhanced output.
Figure 2: SonicMaster dataset creation pipeline and overview.
We introduce a single flow-based generative framework that simultaneously performs dereverberation, equalization, declipping, dynamic-range expansion, and stereo enhancement. The backbone is trained on a curated corpus of polyphonic music rendered through a combinatorial grid of simulated degradations, enabling the network to learn the joint statistics and cross-couplings of common artifacts rather than treating them in isolation. This joint training eliminates the need for error-prone cascades of task-specific modules and reduces inference to a single forward pass.
SonicMaster is conditioned on free-form text prompts that encode high-level production intent. A prompt such as reduce the hollow room sound attenuates late reflections without suppressing desirable early reverberation, whereas increase the brightness selectively enhances the treble frequencies while while preserving spectral balance elsewhere.When the prompt is omitted, an automatic mode leverages learned perceptual heuristics to produce a balanced master suited to non-expert users. Existing models (e.g. VoiceFixer (Liu et al. 2021)) fail on music-focused pipelines typically address artifacts sequentially, ignoring their mutual influence. By unifying restoration and mastering tasks under a single, prompt-driven generative model, SonicMaster delivers professional-grade improvements while affording fine-grained creative control.
In the absence of text-conditioned music-restoration data, we build a new large-scale corpus for controllable restoration. From $\approx 580$ k Jamendo recordings we retain $\approx 25$ k high-quality 30-s segments, balanced across 10 genre groups by production quality score. Each clean clip is corrupted with one to three of 19 common effects drawn from five categories—EQ, dynamics, reverb, amplitude, and stereo—producing paired degraded versions. Every degraded sample is accompanied by a natural-language prompt describing the artifact or required fix, and all random effect parameters are stored as metadata. This genre-diverse collection of tens of thousands of prompt–audio pairs underpins SonicMaster training and offers a rigorous benchmark for controllable music-restoration research. Our main contributions are as follows:
  • We present SonicMaster, a flow-matching model that corrects $19$ common degradations—including reverb, EQ imbalance, clipping, dynamic-range errors, and stereo artifacts in a single generative pass, while allowing free-form textual prompts for fine-grained creative control.
  • To train and evaluate such a system, we curate and release the first text-conditioned music-restoration corpus: 25k high-fidelity Jamendo segments spanning 10 genres, each paired with 7 degraded versions of itself, detailed parameter metadata, and a natural-language instruction describing the required fix, resulting in 175 k audio pairs.
Related Work
Restoring and mastering audio spans speech + music enhancement, audio inpainting, and source separation—areas that have mostly been handled separately (Záviška et al. 2020). Diffusion-based generative models and text-guided audio editing (Hou et al. 2025; Jiang et al. 2025) now let us tackle these problems together. We review these advances, their uses, and the gaps that SonicMaster aims to fill.
Audio Inpainting and Declipping:
Early signal-model and interpolation methods could patch only very short gaps (<10<10 ms), leaving longer dropouts unresolved. Deep generative models now bridge that gap: diffusion-based systems convincingly regenerate missing music sections and clipped peaks (Moliner and Välimäki 2023). The authors in (Wang et al. 2023; Liu et al. 2023) extend this with instruction-guided diffusion for audio inpainting, while VoiceFixer (Liu et al. 2021) jointly denoises, dereverbs, and declipse speech, though it is restricted to voice and does not offer user control. In music, (Imort et al. 2022) removed heavy guitar distortion (including clipping) with neural networks, surpassing sparse-optimization baselines in quality and speed.
Equalization and Tonal Restoration:
Research on learning-based equalization is still emerging. In recently proposed work (Mockenhaupt, Rieber, and Nercessian 2024), the authors introduce CNN based approach to automatically equalize instrument stems by predicting parametric EQ settings, showing improvements over earlier heuristic methods. Notably, the VoiceFixer (Liu et al. 2021) addressed bandwidth extension essentially restoring high-frequency content as part of its speech restoration, which can be seen as a form of equalization correction. Similarly, diffusion-based restorers like MaskSR (Li, Wang, and Liu 2024) treat low-frequency muffling as a distortion to fix, using discrete token prediction to restore a balanced spectrum. In Text2FX (Chu et al. 2025), CLAP (Elizalde et al. 2023) is used in inference mode to steer the parameters of EQ and reverb audio effects.
Method
Dataset
Due to the absence of natural language instruction-prompted music restoration datasets, we have generated SonicMaster music restoration and mastering dataset with text prompts. We source 580580k recordings from the Jamendo [^1] https://www.jamendo.com/ under creative commons licence using the official Jamendo API. To ensure equal representation of genre, we defined $10$ groups, where each group consists of multiple related genre tags, e.g., Hip-Hop genre group containing the following tags: ”rap”, ”hiphop”, ”trap”, ”alternativehiphop”, ”gangstarap”. The full list can be viewed in Appendix. To pick songs for each genre group, we started by deploying the Audiobox Aesthetics toolbox (Tjandra et al. 2025) to assess the production quality of all the recordings. Then, we selected $2,500$ songs for each genre group by filtering songs based on the production quality with an adaptive threshold between 6.586.5-8, depending on the current genre group. We strived to strike a balance between representing all the subtags in each genre group sufficiently enough while maintaining high enough production quality scores. Finally, we extracted a random 30 sec excerpt from each track, positioned between 1515% and 8585% of its total duration. The full data-generation pipeline is illustrated in Figure 2.
We train SonicMaster by applying $19$ distinct degradations to the audio and pairing each with a matching editing instruction. The degradations span five classes: (i) EQ, (ii) Dynamics, (iii) Reverb, (iv) Amplitude, and (v) Stereo.
Equalization (EQ): Spectral degradations cover $10$ effects: Brightness, Darkness, Airiness, Boominess, Muddiness, Warmth, Vocals, Clarity, Microphone, and X-band. Brightness, Darkness, Airiness, Boominess, and Warmth are emulated with low- or high-shelf EQ; Clarity with a Butterworth low-pass filter; Vocals and Muddiness with Chebyshev-II band-pass filters. Microphone applies one of $20$ Poliphone transfer functions (Salvi et al. 2025), while X-band uses an 881212-band, logarithmically spaced peaking EQ with $\textpm 6$ dB gain per band.
Dynamics: This category contains $2$ functions, Compression and Punch: modeled by feedforward compressor and transient shaper, respectively. Both these effects are applied in the time domain with a lossy, irreversible nature, as constructing an exact inverse of these is an ill-posed problem.
Figure 3: Overall architecture of SonicMaster.
Reverb: The Reverb category contains four functions. Three of them utilize the Pyroomacoustics library (Scheibler, Bezzam, and Dokmanić 2018), which simulates acoustic environments with the image source method. We simulate three types of rooms: Small, Big, and Mixed. For our fourth Reverb function, we utilize 12 selected room impulse responses from the openAIR library dataset (Howard and Angus n.d.), which give us audio with more real-life properties. The resulting impulse responses from all the functions are convoluted with the clean signals.
Amplitude: This category consists of two functions: Clipping and Volume. We clip the audio by raising the maximum amplitude to predefined levels, which causes distortion to arise. In Volume, we reduce the volume to a very low (barely audible) range, which decreases the signal-to-quantization noise ratio.
Stereo: This category contains one function to de-stereo the audio recording. First, the track is analyzed for suitability of this operation (whether it is ”stereo” enough) by taking the standard deviation of the difference betwen the left and the right channels and comparing to a set threshold of 0.08, then, if eligible, the left and right channels are combined to create a mono audio recording.
Each ground truth yields $7$ corrupted variants: $4$ with a single, $2$ with double, and $1$ with triple degradation. In multi-degradation, we sample at most one effect from each of the $5$ categories, so an EQ choice, for instance, blocks further EQ picks. To avoid duplicates in the single-degradation set, high-probability effects with narrow parameter ranges: Stereo, Clipping, and Punchare, are used only once.
Each degradation is linked to a one-sentence instruction randomly selected from 8–10 expert-written options; these sentences are concatenated into the full prompt, and we store two prompt variants per clip for robustness. We also record every applied effect and its parameters (gain, absorption), supporting tasks such as parameter prediction. For Compression and Reverb there is a 1515% chance of injecting “hidden clipping” with no corresponding instruction. When neither hidden clipping nor an Amplitude effect is present, the audio is peak-normalised to a random level between 0.81.00.8-1.0. Further details can be found in Appendix.
SonicMaster Architecture
SonicMaster combines Multimodal DiT (Esser et al. 2024) with subsequent DiT blocks (Peebles and Xie 2023). As outlined in Figure 3, the stereo waveform (44.1 kHz) ($x_{t}$) is first encoded by a VAE codec (Evans et al. 2024) into a compact spectro-temporal latent. Restoration therefore occurs entirely in this learned space, allowing large receptive fields without sample-level overhead. The MM-DiT block receives the degraded latent and the text embedding from a frozen FLAN-T5 encoder (Chung et al. 2024), and the conditioned representation is then passed to DiT layers that further process it to predict the velocity $v_{t}$ steering the latent toward its clean target $\hat{x}_{t}$. Prompts like ”reduce reverb” bias this prediction trajectory to suppress decay tails, while the downstream DiT layers also refine musical coherence. A pooled-audio branch, active in 2525% of training cases, concatenates a temporally averaged 5155–15 sec clean cue with the pooled prompt embedding and injects it at every MM-DiT/DiT layer, enabling seamless chaining of $30$ sec segments for long-form generation while degrading gracefully when no reference is supplied.
Audio and text encoding:
We adopt the Stable Audio Open VAE (Evans et al. 2024) to encode–decode stereo signals sampled at 44.1 kHz, yielding a compact latent representation while retaining high-fidelity reconstruction. Text instructions are embedded with FLAN-T5 Large (Chung et al. 2024); the resulting tensor $c_{\text{text}}\in\mathbb{R}^{B\times S_{\text{text}}\times D_{\text{text}}}$ (with $D_{\text{text}}=1024$) is fed to the model as a conditioning signal.
Model (MMDiT/DiT) Clarity Boom Airy Bright Dark Muddy Warm Vocals Microphone X-band
Snippet degraded input
Degraded input 0.0238 0.3601 0.0049 0.0143 0.0893 0.4560 0.4345 0.2525 0.2393 0.1782
Reconstructed input 0.0243 0.3717 0.0051 0.0151 0.0728 0.4749 0.4456 0.2525 0.2379 0.1854
Text2FX-EQ (Chu et al. 2025) 0.0219 0.3809 0.0055 0.0276 0.2112 0.3651 0.4955 0.2199 0.4441 0.3419
SonicMasterSmall (2/6) 0.0100 0.0819 0.0020 0.0064 0.0060 0.0477 0.0590 0.0630 0.0122 0.0408
SonicMasterMedium (4/12) 0.0105 0.0698 0.0021 0.0067 0.0061 0.0400 0.0592 0.0602 0.0091 0.0383
SonicMasterMedium (6/6) 0.0225 0.2766 0.0020 0.0067 0.0056 0.1718 0.1737 0.2417 0.0462 0.0762
SonicMasterLarge (6/18) 0.0114 0.0834 0.0019 0.0059 0.0058 0.0388 0.0617 0.0576 0.0088 0.0358
   -w 5sec Audio Cond. 0.0111 0.0716 0.0021 0.0061 0.0058 0.0386 0.0605 0.0628 0.0124 0.0387
   -w 15sec Audio Cond. 0.0117 0.0750 0.0020 0.0064 0.0063 0.0320 0.0552 0.0525 0.0079 0.0398
   -w/o Audio Cond. (basic) 0.0099 0.0658 0.0021 0.0064 0.0056 0.0352 0.0595 0.0746 0.0097 0.0434
   -w Cond. During Infer 0.0115 0.0840 0.0019 0.0060 0.0058 0.0389 0.0610 0.0572 0.0088 0.0355
   -w/o Text Cond. 0.0130 0.1432 0.0032 0.0101 0.0086 0.0448 0.0841 0.0668 0.0154 0.0424
   -w Euler $1$ Step 0.0100 0.1146 0.0019 0.0059 0.0061 0.0425 0.0668 0.0498 0.0141 0.0384
   -w Euler $100$ Steps 0.0136 0.1540 0.0033 0.0100 0.0091 0.0540 0.0915 0.0749 0.0162 0.0444
   -w Runge-Kutta $10$ Steps 0.0120 0.0810 0.0019 0.0058 0.0058 0.0402 0.0630 0.0590 0.0083 0.0374
Fullsong degraded input
Degraded Input 0.0290 0.3231 0.0048 0.0124 0.0983 0.4606 0.4810 0.2274 0.2403 0.1737
SonicMasterLarge (6/18) 0.0102 0.0639 0.0021 0.0060 0.0065 0.0329 0.0510 0.0517 0.0070 0.0289
Table 1: Objective evaluation of EQ attribute improvement with average absolute error values – the lower, the better.
Model (MMDiT/DiT) Reverb Dynamics Amplitude Stereo
Small Big Mix Real Compressor Punch Clip Volume
Snippet degraded input
Degraded input 0.4457 0.4243 0.5045 0.4639 0.0496 0.1200 5.122 0.1813 0.4183
Reconstructed input 0.4686 0.4507 0.5433 0.4908 0.0494 0.0590 3.871 0.1810 0.4181
HPSS 6 dB 0.4419 0.4240 0.4970 0.4537 - - - - -
HPSS 12 dB 0.4971 0.4739 0.5333 0.4814 - - - - -
WPE (Nakatani et al. 2010) 0.4849 0.4732 0.5207 0.4854 - - - - -
SonicMasterSmall (2/6) 0.3812 0.3826 0.4050 0.3277 0.0172 0.0859 2.363 0.0457 0.1536
SonicMasterMedium (4/12) 0.3683 0.3700 0.3934 0.3138 0.0147 0.0891 2.455 0.0409 0.1028
SonicMasterMedium (6/6) 0.3952 0.3916 0.4422 0.4255 0.0366 0.0833 2.905 0.1228 0.4180
SonicMasterLarge (6/18) 0.3663 0.3726 0.3935 0.3109 0.0193 0.0871 1.506 0.0468 0.1058
   -w 5sec Audio Cond. 0.3717 0.3658 0.3919 0.3079 0.0164 0.0893 1.779 0.0430 0.0918
   -w 15sec Audio Cond. 0.3676 0.3682 0.3901 0.3093 0.0172 0.0895 1.633 0.0485 0.1008
   -w/o Audio Cond. During Training 0.3620 0.3682 0.3888 0.3067 0.0146 0.0850 2.055 0.0455 0.1015
   -w Cond. During inference 0.3664 0.3724 0.3934 0.3112 0.0172 0.0870 1.455 0.0412 0.1060
   -w/o Text Cond. 0.3732 0.3805 0.4012 0.3264 0.0157 0.0730 2.812 0.0465 0.1416
   -w Euler 1 step 0.4215 0.4378 0.4599 0.3459 0.0124 0.0906 2.171 0.0461 0.1261
   -w Euler 100 Steps 0.3716 0.3754 0.3997 0.3255 0.0158 0.0672 2.753 0.0491 0.1497
   -w Runge-Kutta 10 Steps 0.3647 0.3684 0.3921 0.3087 0.0210 0.0858 1.422 0.0481 0.1059
Full song degraded input
Degraded input 0.3667 0.3654 0.4706 0.3852 0.0598 0.1103 6.363 0.1829 0.4133
SonicMasterLarge (6/18) 0.3954 0.4511 0.4191 0.4066 0.0258 0.1101 3.734 0.0424 0.0850
Table 2: Objective evaluation: Reverb, Dynamics, Amplitude, and Stereo improvement. Clip values are multiplied by 1000.
Category Text relevance Quality1 Quality2 Consistency Preference
EQ 5.11 $\pm$ 0.55 4.55 $\pm$ 0.44 5.01 $\pm$ 0.52 5.05 $\pm$ 0.42 4.84 $\pm$ 0.49
Reverb 5.59 $\pm$ 0.57 4.25 $\pm$ 0.68 5.28 $\pm$ 0.57 5.31 $\pm$ 0.54 5.27 $\pm$ 0.75
Dynamics 4.56 $\pm$ 0.80 4.09 $\pm$ 0.54 5.00 $\pm$ 0.81 5.00 $\pm$ 0.84 4.91 $\pm$ 1.01
Amplitude 6.19 $\pm$ 0.66 3.66 $\pm$ 0.43 5.38 $\pm$ 0.77 5.59 $\pm$ 0.81 5.75 $\pm$ 0.84
Stereo 5.50 $\pm$ 1.05 4.62 $\pm$ 0.83 5.50 $\pm$ 0.63 5.38 $\pm$ 0.70 5.25 $\pm$ 0.77
Mixed degradations 5.28 $\pm$ 0.51 3.85 $\pm$ 0.88 4.60 $\pm$ 0.59 4.65 $\pm$ 0.62 4.93 $\pm$ 0.57
Table 3: Listening study results - Mean Opinion Score with 95% confidence interval.
Model Single deg. Double+triple deg. All
FAD $\downarrow$ KL $\downarrow$ SSIM $\uparrow$ PQ $\uparrow$ FAD $\downarrow$ KL $\downarrow$ SSIM $\uparrow$ PQ $\uparrow$ FAD $\downarrow$ KL $\downarrow$ SSIM $\uparrow$ PQ $\uparrow$
Snippet degraded input
Ground truth mastered reference - - - 7.886 - - - 7.886 - - - 7.886
Degraded input 0.061 3.859 0.838 7.321 0.184 6.827 0.696 6.632 0.106 5.131 0.777 7.026
Reconstructed input 0.139 3.990 0.574 7.172 0.290 6.984 0.507 6.501 0.196 5.273 0.546 6.885
SonicMasterSmall (2/6) 0.071 0.726 0.623 7.716 0.088 1.215 0.586 7.609 0.077 0.935 0.607 7.670
SonicMasterMedium (4/12) 0.070 0.709 0.624 7.740 0.084 1.187 0.589 7.649 0.075 0.914 0.609 7.701
SonicMasterMedium (6/6) 0.086 1.893 0.603 7.571 0.154 3.241 0.555 7.231 0.110 2.470 0.583 7.426
SonicMasterLarge (6/18) 0.069 0.696 0.624 7.743 0.082 1.145 0.589 7.654 0.073 0.888 0.609 7.705
   -w 5sec Audio Cond. 0.070 0.703 0.624 7.733 0.083 1.175 0.588 7.637 0.075 0.905 0.609 7.692
   -w 15sec Audio Cond. 0.069 0.694 0.623 7.742 0.083 1.161 0.588 7.650 0.073 0.894 0.608 7.702
   -w/o Audio Cond. During Training 0.069 0.691 0.625 7.741 0.082 1.146 0.590 7.645 0.073 0.886 0.610 7.700
   -w Cond. During Inference 0.069 0.693 0.625 7.742 0.082 1.141 0.589 7.653 0.073 0.885 0.609 7.704
   -w/o Text Cond. 0.069 0.917 0.621 7.772 0.088 1.484 0.586 7.643 0.074 1.160 0.606 7.716
   -w Euler 1 step 0.076 0.922 0.615 7.684 0.117 1.789 0.567 7.520 0.090 1.294 0.594 7.614
   -w Euler 100 Steps 0.069 0.920 0.620 7.764 0.087 1.521 0.585 7.621 0.076 1.178 0.605 7.703
   -w Runge-Kutta 10 Steps 0.070 0.701 0.624 7.740 0.084 1.171 0.588 7.642 0.074 0.902 0.608 7.698
Full song degraded input
Ground truth mastered reference - - - 7.885 - - - 7.885 - - - 7.885
Degraded input 0.087 2.937 0.834 7.325 0.223 5.679 0.682 6.606 0.142 4.308 0.758 6.965
Reconstructed input 0.165 3.049 0.584 7.204 0.335 5.644 0.510 6.509 0.234 4.339 0.547 6.859
SonicMasterLarge (6/18) 0.095 0.754 0.380 7.627 0.121 1.251 0.368 7.477 0.101 1.002 0.374 7.552
Table 4: Objective evaluation with FAD, KL, SSIM, and PQ metrics. For readability, KL values were multiplied by 1000.
Training with the rectified flow:
We train SonicMaster using rectified flow (Liu, Gong, and Liu 2022; Esser et al. 2024), to predict the flow velocity from degraded to clean audio in latent space, unlike other models that map noise to output distributions (Fei et al. 2024; Hung et al. 2024).
We assign timestep $t=1$ to the latent representation of the degraded audio $x_{1}$, and $t=0$ to the latent representation of the clean audio target $x_{0}$. During training, we feed the model with samples $x_{t}$, which are interpolations of degraded input $x_{1}$ and clean target $x_{0}$ on a straight line:
$x_{t}=tx_{1}+(1-t)x_{0}$ (1)
where timestep $t$ is drawn from a skewed distribution with increasing probability for higher $t$ as follows:
$p(t)=0.5U(t)+0.5t,\quad t\in[0,1]$ (2)
where $U$ represents a uniform distribution. This skewed distribution gives emphasis to more degraded inputs given the interpolation of training data in Eq. 1.
The model is trained to predict the flow velocity $v_{t}$ from the current $x_{t}$ to the target clean audio $x_{0}$:
$v_{t}=\frac{dx_{t}}{dt}=x_{0}-x_{1}$ (3)
The model $f_{\theta}$ with parameters $\theta$ estimates the velocity $\hat{v_{t}}$:
$f_{\theta}(x_{t},t,c_{text})=\hat{v_{t}}$ (4)
where $c_{text}$ is the text condition from the FLAN-T5 model, which is passed to the dual-stream MM-DiT blocks as one of the streams. The $c_{text}$ condition is also passed through a pooled projection and used to control the scale and shift factors of the adaptive layer-norm layers in both MM-DiT and DiT blocks. The training loss is then given as:
[…]
During inference, the degraded audio input $x_{1}$ is converted into a clean audio output $x_{0}$ by integrating the predicted velocity $\hat{v_{t}}$ using the forward Euler method:
$x_{t+h}=x_{t}+h\hat{v_{t}}$ (6)
where $h\in[0,1]$ is the step computed as the inverse of total timesteps dedicated for integration.
Inference
During inference, SonicMaster takes in an audio input and a text instruction given by user to perform desired restoration/mastering operation. Inference is possible without text input in the so-called auto-correction mode. To process full-length songs, SonicMaster operates on chunks of 30 sec and then connects the segments together. After the first segment is inferred, the last 10 sec of this output are used to condition the next segment inference through the audio pooling branch. The overlapping regions of the resulting segments are then linearly interpolated over the overlapping 10 seconds to connect the segments together.
Experimental Setup and Baselines
Baselines and training setup
We train SonicMaster using NVIDIA L40S GPUs for 40 epochs. We adopt classifier-free guidance (Ho and Salimans 2022) by (i) dropping the text prompt in 1010% of samples and (ii) replacing it in another 1010% with one of four generic phrases (“Make it sound better!”, “Master this track for me, please!”, “Improve this!”, “Can you improve the sound of this song?”). In 2525% of cases, the model is additionally conditioned—via the pooling branch—on the first 10 s of clean audio. Unless stated otherwise, all experiments follow these conditioning settings while comparing multiple SonicMaster variants and baselines.
[…]
Objective Evaluation
Evaluation is conducted along two orthogonal axes. (i) Global perceptual fidelity is quantified with Frechét Audio Distance (FAD) on CLAP embeddings (Elizalde et al. 2023), Kullback–Leibler divergence (KL), structural similarity (SSIM) on 128-bin mel-spectrograms, and the Production Quality (PQ) score from the Audiobox Aesthetics toolbox (Tjandra et al. 2025). (ii) Degradation-specific restoration efficacy is measured by average absolute error reduction: for every degraded clip in a $7000$ clip test set, we compute the relevant (based on the degradation deployed) artefact-aware metric against its clean counterpart from a $1000$ sample reference set, then recompute the metric after SonicMaster processing; the relative decrease indicates how closely each model variant approaches the ground-truth.
For X-band EQ and microphone-TF degradations we compute the spectral balance over eight bands and report their cosine distance. All other EQ effects are scored by the energy ratio between the affected band and the full spectrum. Compression is measured as the standard deviation of frame-level RMS (2048-sample frames, 1024 hop); punch as the mean onset-envelope value (librosa.onset.onset_strength). Because RT60 estimates are unreliable on dense mixes, reverb is assessed via the Euclidean distance of modulation spectra. Clipping uses spectral flatness; volume, the global RMS; and stereo width, the RMS ratio of the mid and side signals, $\mathrm{RMS}\!\left[\frac{L-R}{2}\right]/\mathrm{RMS}\!\left[\frac{L+R}{2}\right]$.
Subjective Evaluation
We presented listeners with $43$ audio sample pairs – degraded inputs and SonicMaster outputs – to rate, consisting of $2$ pairs for each degradation function ($2\times 19=38$ single degraded samples), $3$ pairs of double and $2$ pairs of triple degraded samples. Using a $7$-point Likert Scale, listeners were to rate: 1) The extent of improvement from the input to SonicMaster output represented by the text prompt (Text relevance), 2) audio quality of input (Quality1), 3) audio quality of the inferred SonicMaster sample (Quality2), 4) consistency and fluency of the inferred sample (Consistency), and 5) preference between the two samples, where 1 represents full preference of the ground truth degraded input, and 7 represents the SonicMaster inferred sample (Preference). The study was attended by 8 listeners (5 music experts and 3 Music Information Retrieval researchers).
(a) Ground truth audio.
(b) Degraded by reverb (smearing).
(c) After SonicMaster: reverb removed.
Figure 4: Comparison of spectrograms: (a) ground truth, (b) degraded with reverb, and (c) the output of SonicMaster where smearing is removed. Prompt: ”Please, reduce the strong echo in this song.”
Results
The results of the attribute controllability evaluation are depicted in Tables 1 and 2. SonicMaster significantly outperforms the baselines of Text2FX in EQ, and WPE and HPSS in Reverb. SonicMaster shows a clear improvement in all categories when compared to the degraded and reconstructed inputs. Furthermore, we can observe that the reconstructed input metrics are overall slightly worse (with exceptions) than those of the ground truth degraded inputs.
Across the audio-quality metrics in Table 4, SonicMaster outperforms the degraded inputs in both PQ and KL. Its FAD is only marginally higher than that of the degraded audio, yet markedly lower than the reconstructed baseline. Furthermore, SonicMaster achieves a significant increase in PQ, almost reaching the level of ground truth mastered reference. In SSIM, SonicMaster exhibits lower scores than the degraded inputs but achieves superior performance compared to the reconstruction baseline. From Tables 1, 2, 4, we can observe that SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. SonicMasterMedium (4/12) performs slightly better than the SonicMasterSmall, but still lacks behind SonicMasterLarge in Clip. SonicMasterMedium (6/6) performs the worst out of all variants across all metrics.
As part of our ablation study, we compared SonicMasterLarge with different settings. Models with different lengths of audio conditioning (5 sec, 15 sec) perform similarly to the 10 sec setting. We chose to use 10 sec setting as our default as a compromise between computation speed and length of overlap in neighbouring segments. Further, conditioning during inference, or no conditioning during training and inference at all produced comparable results. While the no condition variant provides best performance in Boom with absolute average error of 0.0658, it also provides the worst performance in Clip error with 2.055 compared to the aforementioned variants. Furthermore, we evaluated inference without text prompts, which yields comparable performance in FAD, SSIM, and PQ, but shows worse performance in KL (0.917 vs 0.696 on single degradation snippets). Performance in Clip and Stereo is also worse with values of 2.812 and 0.1416 as opposed to 1.506 and 0.1058. In EQ, values across the error values across the board are slightly higher for this variant, as the model is not targetting a specific EQ enhancement due to the missing text prompt.
We evaluated Euler solvers with $1$, $10$ (baseline), and $100$ steps, plus a $10$-step 4th order Runge–Kutta (Dormand and Prince 1980) solver. Euler-1 matches the baseline overall but is weaker on Boom, Microphone, Clip, all Reverb subtasks, and shows higher KL. Euler-100 boosts Reverb and Punch yet lowers every EQ score versus the 1-/10-step runs. Runge–Kutta-10 equals Euler-10 on most metrics and tops Clip, but its inference is significantly slower.
When evaluating SonicMaster with full songs, we can again observe significant improvements in EQ metrics (Table 1), and in most functions in Table 2. However, the reverb metric of MSD shows mixed results. In audio quality metrics, SonicMaster shows a decrease in SSIM and FAD compared to degraded inputs, except for FAD in double and triple degradation samples, which points to SonicMaster’s ability to handle multiple degradations at once.
Subjective evaluation results are depicted in Table 3. In text relevance, we can observe very good ratings in Amplitude (6.19) and Reverb categories (5.59), which highlight significant perceptive improvement in declipping, volume increase, and in dereverberation. These two categories further show the highest improvement in quality, highest consistency ratings, and highest preference. EQ shows the third best text relevance, but worst preference ratings. Overall, SonicMaster samples are rated higher in quality compared to inputs and preferred across the board. A paired t-test on Quality1 and Quality2 ratings shows statistically significant differences in Reverb, Amplitude, Stereo, and mixed degradations. These results further confirm SonicMaster’s capabilities in text-controllable music restoration and mastering.
Discussion
The experiments confirm that SonicMaster improves audio quality, proving that a generative strategy for music restoration and mastering is feasible when backed by a large corpus and a suitable training objective. A key limitation, however, are occasional artefacts introduced by the lossy latent representation: reconstructed clips can exhibit “robotic” vocals or muted instruments, especially for certain genres, and the effect is stronger for degraded inputs than for their clean references. Future research should pursue less lossy latent encoders or alternative representations to minimize these distortions.
The observed decrease in SonicMaster’s performance on full songs in SSIM and Reverb metrics could be related to the way neighbouring segments are connected together. Improving on this aspect could increase the objective performance further. Moreover, evaluating the amount of reverberation in music is challenging, as the audio is usually very dense. Additionally, the way a generative model such as SonicMaster removes reverb in the latent space is not telegraphed, which makes it more difficult to choose a correct metric. A deep exploration of this topic would be of use to the community.
Conclusion
We introduced SonicMaster, the first unified text-guided generative model for music restoration and mastering, capable of addressing a wide range of audio degradations within a single model. Key contributions include the creation of a new SonicMaster dataset [^2] Dataset, code, samples, model, and demo available through https://amaai-lab.github.io/SonicMaster/ of paired degraded and high-quality music with textual annotations and the use of a flow-matching training paradigm to directly learn the restoration mapping. We also developed a text-conditioned multimodal generative architecture that unifies diverse enhancement tasks, enabling the model to perform all restoration and mastering steps in one generative pass. Comprehensive evaluations demonstrate that SonicMaster substantially improves audio quality, achieving strong results in objective metrics (e.g., FAD, KL divergence, SSIM, PQ) and consistently earning higher listener preferences in blind tests compared to both original and baseline outputs.
Appendix A Appendix
Genre tags
We grouped genre tags into genre groups, as depicted in Table 5. Each row links a coarse “Group” label—such as Rock, Electronic, or Jazz/Blues—to the fine-grained “Genre tags” that appear in the metadata. These tags enumerate substyles (e.g., progressiverock, deephouse, acidjazz), which allows us to aggregate diverse representations inside each of the genre grops.
Table 5: Genre groupings by metadata tags used in our dataset.
Group Genre tags
Rock rock, alternativerock, poprock, classicrock, hardrock, progressiverock, stoner, psychedelicrock, garage, indierock
Pop pop, electropop, dancepop, dance, alternativepop, adultcontemporary, indiepop
Electronic electronic, house, techno, trance, edm, electrohouse, deephouse, progressivehouse, electroswing, synthwave, electronica
Hip-Hop rap, hiphop, trap, alternativehiphop, gangstarap
Folk folk, singersongwriter, americana, country, bluegrass, folklore
Metal metal, deathmetal, blackmetal, thrashmetal, heavymetal, numetal, metalcore, hardcore, alternativemetal, doommetal
World world, latin, reggaeton, afrobeat, african, indian, oriental, celtic, salsa, flamenco, jpop, middleeastern, asian, reggae
Jazz/Blues jazz, blues, funk, acidjazz, jazzfusion, smoothjazz, jazzfunk, soul, swing, rnb, alternativernb
Chill ambient, downtempo, chillout, chillhop, lofi, newage, darkambient, triphop, chillwave, idm, dreampop
Classical classical, filmscore, neoclassical, symphonic, opera, baroque, medieval, avantgarde, production, choral
Degradation functions
To create the SonicMaster dataset, we used a set of 19 degradation functions. The details of their implementation and parameter range are described in Table 6. Each of the groups, and subsequently each of the functions inside the groups, have their own probabilities/weights to be picked in our data creation pipeline. These are documented in Table 7.
Table 6: Detailed description of degradation functions used to create our dataset.
Degradation group Degradation type Description Prompt example (inverse)
EQ X-band EQ Apply 8 to 12 band parametric EQ with 6-6 to +6+6 range for each band. Correct the unnatural frequency emphasis.
Microphone transfer function Convolve the audio with one of 20 phone microphone transfer functions. Reduce the coloration added by the microphone.
Brightness Reduce brightness using a high-shelf filter at 6 kHz by 6–15 dB. Give the mix more shine and sparkle.
Darkness Increase perceived brightness with a high-shelf filter at 6 kHz by 6–15 dB. Make the tone fuller and less sharp.
Airiness Reduce airiness via a high-shelf filter at 10 kHz by 10–20 dB. Add more air and openness to the sound.
Boominess Reduce boominess with a low-shelf filter at 120 Hz by 10–20 dB. Give the audio more roar and low-end power.
Clarity Degrade clarity using a Butterworth low-pass filter (order 3–5) with cutoff at 2 kHz. Increase the clarity of this song by emphasizing treble frequencies.
Muddiness Increase muddiness with a 2nd-order Chebyshev Type II bandpass (200–500 Hz) by 6–15 dB. Make the mix sound less boxy and congested.
Warmth Reduce warmth with a low-shelf filter at 400 Hz by 6–20 dB. Make the sound warmer and more inviting.
Vocals Attenuate vocal-range frequencies using a 2nd-order Chebyshev Type II bandpass (350–3500 Hz) by 6–20 dB. Make the vocals stand out more.
Dynamics Compression Apply a feedforward compressor with attack 3–80 ms, release 80–250 ms, threshold 45-45 to 38-38 dB, ratio 6–45, and make-up gain 16–25 dB. Let the audio breathe more and improve the dynamics.
Punch Apply a feedforward transient shaper with attack 3 ms, release 150 ms, adaptive threshold, and reduction of 8–15 dB. Add more impact and dynamic punch to the sound.
Reverb Small room Convolve with Pyroomacoustics simulated IR: room size $(7\text{--}15,8\text{--}18,4\text{--}14)$ m, absorption coefficient 0.05–0.30. Clean this off any echoes!
Big room Convolve with Pyroomacoustics IR: room size $(4\text{--}8,4\text{--}7,2.5\text{--}3.5)$ m, 1–2 absorptive walls, frequency-dependent absorption. Can you remove the excess reverb in this audio, please?
Mixed material room Convolve with Pyroomacoustics IR: room size $(3\text{--}7,3\text{--}9,2.5\text{--}4)$ m, absorption coefficient 0.05–0.30. Remove excess reverb and make it sound cleaner.
Real RIR Apply one of twelve real impulse responses from the openAIR library. Please, reduce the strong echo in this song.
Amplitude Clipping Modify the audio level to a maximum amplitude of {2,3,5} and apply clipping. Reduce the clipping and reconstruct the lost audio, please.
Volume Adjust the audio gain to a maximum amplitude of {0.001, 0.003, 0.01, 0.05}. Enhance the loudness without distorting the signal.
Stereo Stereo Combine the left and right channels to erase the spatial image. Add depth and separation between left and right.
Group (weight) Option Probability / Weight
EQ (0.4) xband 7.0
mic 5.0
bright 3.0
dark 3.0
airy 2.0
boom 2.0
clarity 3.0
mud 3.0
warm 3.0
vocal 4.0
Dynamics (0.125) comp 2.5
punch 1.0
Reverb (0.225) small 0.15
big 0.15
mix 0.30
real 0.40
Amplitude (0.125) clip 3.0
volume 1.0
Stereo (0.125) stereo 1.0
Table 7: Degradation groups with assigned probabilities and option weights.
(a) Original (clean).
(b) Degraded input.
(c) Restored output.
Figure 5: Original vs. degraded (via convolution with a phone microphone transfer function) and SonicMaster-restored spectrograms; restoration suppresses the microphone’s coloration.
Spectrogram examples
We visualize time–frequency structure in spectrograms to provide qualitative evidence of restoration behavior. Each figure shows the clean reference, the degraded input (e.g., reverberation-induced smearing or clipping distortion), and the output of SonicMaster. Figures 5, 6, 7, 8, and 9 compare clean, degraded, and restored spectrograms across selected scenarios (reverb, clipping, microphone transfer function, and clarity EQ).
(a) Original (clean).
(b) Degraded input.
(c) Restored output.
Figure 6: Effect of reverberation (example from the main text in larger size): top panel shows the original audio sample, middle panel shows audio convolved with a Pyroomacoustics simulated impulse response, and bottom panel shows the dereverberated result with echoes cleaned.
(a) Original (clean).
(b) Degraded input.
(c) Restored output.
Figure 7: Effect of clarity degradation and restoration on spectrograms. The treble frequencies are supressed in the degraded input sample, and then restored with SonicMaster. Prompt: ”Make the audio clearer and more intelligible.”
(a) Original (clean).
(b) Degraded input.
(c) Restored output.
Figure 8: Effect of clipping degradation and related restoration. Drum hits clip in the degraded audio, but are restored in the SonicMaster’s output without distortion. Prompt: ”Clean up the harshness in the signal.”
(a) Original (clean).
(b) Degraded input.
(c) Restored output.
Figure 9: Another example of the effect of clipping and its restoration. The degraded input shows signs of distortion with visible spectral content above 15 kHz. This distortion is suppressed by SonicMaster. Prompt: ”Clean up the noisiness in the audio.”
Prompts for each degradation type
Prompt instructions for each degradation type are grouped by audio attribute in Table 8; for example, entries for Xband, microphone coloration, clarity, brightness, darkness, airiness, boominess, warmth, muddiness, vocals, compression, punch, reverb, volume, clipping, and stereo give natural-language commands that steer the restoration model. These instructions act as conditioning signals—e.g., “remove excess reverb and make it sound cleaner,” “raise the level of the vocals,” or “make this sound brighter”—so that the generative restoration trajectory emphasizes or suppresses specific signal characteristics.
Table 8: User instructions grouped by audio attribute.
Attribute Example Instructions
Xband Can you please correct the equalization?; Improve the balance in the audio by fixing the chaotic equalizer, please.; Make this sound balanced, please.; Balance the EQ, please.; Balance the tonal spectrum of the audio.; Correct the unnatural frequency emphasis.; Make the EQ curve smoother and more natural.; Even out the EQ.; Adjust the tonal balance for a more pleasing sound.
Microphone This audio was recorded with a phone, can you fix that, please?; Please make this sound better than a phone recording.; Balance the EQ, please.; Improve the balance in this song.; Make the audio sound like it was recorded with a higher-quality microphone.; Reduce the coloration added by the microphone.; Make the tone more neutral and balanced.; Improve the naturalness of the recording.; Remove the harshness or boxiness from the mic coloration.
Clarity Increase the clarity!; Can you please make this song sound more clear?; Increase the clarity of this song by emphasizing treble frequencies.; Make the audio clearer and more intelligible.; Sharpen the overall sound.; Bring more focus and definition to the details.; Make the mix sound less cloudy.; Tighten the articulation in the sound.
Brightness Can you please make this sound brighter?; Increase the brightness!; Make this audio sound brighter by emphasizing the high frequencies.; Add some brightness to the high end.; Make the sound more vivid and lively.; Give the mix more shine and sparkle.; Lift the treble for a more open tone.; Enhance the presence of the upper frequencies.
Darkness Make this sound darker!; Can you reduce the brightness, please?; Make the audio darker by suppressing the higher frequencies.; Bring in more low-mid richness to make the sound darker.; Make the tone fuller and less sharp.; Smooth out the highs with deeper low-end support.; Round out the sound with more body.; Soften the harshness with a warmer tone.
Airiness Make this sound more fresh and airy by emphasizing the high end frequencies.; Make this feel more airy, please.; Increase the perceived airiness, please.; Give this a light sense of spaciousness by amplifying the higher frequencies.; Add more air and openness to the sound.; Make the audio feel more spacious and extended.; Enhance the sense of space in the highs.; Lift the top end for a more open character.; Give the mix a breathier, more open feel.
Boominess Make it boom!; Make this song sound more boomy by amplifying the low end bass frequencies.; Increase the boominess, please!; Give me more bass!; Can you make this more bassy, please?; Give the audio more roar and low-end power.; Make the bass more impactful and solid.; Add weight and depth to the bottom end.; Reinforce the low frequencies for more energy.; Boost the bass presence.
Warmth Can you make this song sound warmer, please?; Increase the warmth, please.; Emphasize the bass and low-mid frequencies to give this a more warm feel.; Make the sound warmer and more inviting.; Add some low-mid warmth to the mix.; Soften the tone with a bit more body.; Give the audio a warm analog feel.; Enhance the warmth for a fuller sound.
Muddiness Can you make this song sound less muddy, please?; Decrease the muddiness!; Reduce the level of muddiness in this audio by lowering the low-mid frequencies.; Clean up the muddiness in the low-mids.; Make the mix sound less boxy and congested.; Improve definition by reducing mud.; Clear up the low-mid buildup.; Make the audio tighter and less murky.
Vocals Raise the level of the vocals, please.; Can you amplify the vocals, please?; Emphasize the vocals by raising the level of the mid frequencies specific for vocals.; Bring the vocals forward in the mix.; Make the voice clearer and more present.; Increase the vocal presence by enhancing the midrange.; Make the vocals stand out more.; Strengthen the vocal clarity and focus.
Compression Increase the dynamic range.; Decompress the audio, please.; Remove the compression, please.; Can you fix the strong compression effect in this audio by expanding the dynamic range?; Restore the dynamics of the audio.; Make the sound less squashed and more open.; Reduce the over-compression for a more natural feel.; Bring back the contrast in volume.; Let the audio breathe more and improve the dynamics.
Punch Give this song a punch!; Make the transients sharper, please.; Increase the punchiness of the song by emphasizing the transients.; Make the audio more punchy and energetic.; Bring back the snap and attack of transients.; Add more impact and dynamic punch to the sound.; Make drums and hits sound more aggressive and tight.; Increase the percussive clarity and definition.
Reverb Can you remove the excess reverb in this audio, please?; Please, dereverb this audio.; Remove the echo!; Please, reduce the strong echo in this song.; Remove the church effect, please.; Clean this off any echoes!; This song has too much reverb present, can you reduce it?; Make the audio sound more dry and direct.; Reduce the roominess or echo.; Remove excess reverb and make it sound cleaner.; Bring the sound closer and more focused.; Tighten the spatial feel of the audio.
Volume The volume is low, make this louder please!; Can you make this sound louder, please?; Increase the amplitude.; Normalize the audio volume.; Make the audio louder and more powerful.; Increase the overall level.; Boost the volume without distorting the signal.
Clipping This audio is clipping, can you please remove it?; Remove the loud hissing in this song?; Remove the clipping.; Reduce the clipping and reconstruct lost audio.; Clean up noisiness.; Make the audio smoother and less distorted.; Reduce the gritty or crushed character.; Fix digital distortion.
Stereo Make it sound spacious!; Can you make this audio stereo, please?; Alter left/right channels to give spatial feel.; Widen the stereo image.; Add depth and separation between left and right.; Enhance the stereo field for immersive sound.